what is the pH of 0.36M methylamine solution?

pH of 0.36 M methylamine sol. (Kb = 4.4 * 10^-4)
Answer #1

CH3NH2 (aq)+ H2O(l) --> CH3NH3+(aq) + OH- (aq)

Kb= [CH3NH3+] [OH-]/ [CH3NH2]

If X mole of methylamine ionised.

Then

Kb = X × X / .36-X

4 × 10-4= X​​​​​​2/.36 (X<<<<.36)

X​​​​​​​​​​​2 = 4.4 ×10-4× .36

X = 1.2586×10-2

X= .012586

pOH = - log [OH-]

= - log X= - log .012586

= 1.9

pH = 14-pOH = 14- 1.9

= 12.1